please answer each part! ill give positive feedback!!!


media%2F391%2F39132ebf-9d0b-43b0-b37a-38please answer each part! ill give positive feedback!!!
Answer #1

a.

Value of each Debenture = 1,620,000/1620

Value of Each Debenture = $1,000

Coupon Rate = 7%

Dollar amount of interest per bond per year = 0.07(1000)

Dollar amount of interest per bond per year = $70

b.

Total Interest Expense per year = 70*1620

Total Interest Expense per year = $113,400

c.

After tax interest cost = (1 - 0.35)11340

After tax interest cost = $73,710